OpenStudy (anonymous):

Find Derivative: 3xy^2 + y^3 = 8

OpenStudy (anonymous):

with respect to x or y?

OpenStudy (anonymous):

derivative with respect to x I am assuming

OpenStudy (anonymous):

im not sure says to find in terms of x and y using implicit

OpenStudy (anonymous):

now just solve for y'

OpenStudy (anonymous):

i got y/2x+y is that correct?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Hold on ,

OpenStudy (anonymous):

3xy^2 + y^3 = 8 using product rule 3(2y y' x + y^2) +3y^2 y'=0

OpenStudy (anonymous):

now just solve for y'

OpenStudy (anonymous):

ohhh with the 3 as a constant I have to put it outside the brackets of the whole derivative of the 3xy^2?

OpenStudy (anonymous):

no, just one bracket 3 multiplies into just one bracket 3(2y y' x + y^2)

OpenStudy (anonymous):

right I understand. so is my final answer of y/2x+y correct?

OpenStudy (anonymous):

Your answer was very nearly correct amzerz

OpenStudy (anonymous):

You are missing a negative I think.

OpenStudy (anonymous):

minus

OpenStudy (anonymous):

y' = -y/(2x+y)

OpenStudy (anonymous):

right I see that now! thanks!!